# Do Robots Really Need Anthropomorphic Hands?

### v1

@@ Abstract
Human manipulation skills represent a pinnacle of their voluntary motor functions, requiring the coordination of many degrees of freedom and processing of high-dimensional sensor input to achieve such a high level of dexterity. Thus, we set out to answer whether the human hand, with its associated biomechanical properties, sensors, and control mechanisms, is an ideal that we should strive for in robotics—do we really need anthropomorphic robotic hands?
@@
Although complex five-fingered hands are often used as the ultimate goal for robotic manipulators, they are not necessary for all tasks. We found that wrist flexibility and finger abduction/adduction are important for manipulation capabilities. On the contrary, increasing the number of fingers, actuators, or degrees of freedom is often not necessary. Three fingers are a good compromise between simplicity and dexterity. Non-anthropomorphic hand designs with two opposing pairs of fingers or human hands with six fingers can further increase dexterity, suggesting that the human hand may not be the optimum.
@@
Robotics Hands, Grippers, Smart Prosthetics, Manipulation, Grasping
@@ 1 Introduction
All of which prompts the question of whether we need anthropomorphic robotic hands. Attempting to answer this question, we analyze the human hand and its capabilities, including kinematics and perception, to find characteristic features of anthropomorphic hands. Furthermore, we review commercially available robotic and prosthetic hands to identify anthropomorphic features that are technically feasible. Then we perform a systematic literature review to correlate hand features derived from the previous two sections, skill categories, and manipulable degrees of freedom. We conclude with a discussion that connects our findings to literature from neuroscience, biology, prosthetics, and robotics.
Figure 1: Kinematics of the human hand (legend in Figure 7(a)).
@@ 2 The Human Hand
2.1 Biomechanics of the Human Hand
The human hand is one of the most advanced manipulation devices that can be found in nature [41].
@@ 2.1.1 Joints and Degrees of Freedom
The human hand has many bones (see Figure 2), but not all of them are connected through movable joints. The human hand has 20 undeniable degrees of freedom—four in each finger. Index, middle, ring, and little finger can perform flexion and extension in three joints (metacarpo-phalangeal joint, MCP; proximal interphalangeal joint, PIP; distal interphalangeal joint, DIP) and adduction (moving closer to the middle finger) and abduction (moving away from the middle finger) in one joint (MCP). The thumb is a special case with a saddle joint (trapezo-metacarpal joint, TMC) that has two degrees of freedom to achieve flexion/extension and opposition/reposition, and two more joints (MCP and interphalangeal joint, IP) to perform flexion/extension. It is possible to move the metacarpal bones and hollow the palm to a limited extent [41] so that [15] define hand kinematics with 24 degree of freedom. Assuming the carpus to be rigid is only an approximation [41], and the palm, as well as the skin, are soft and deformable. A kinematic diagram of the hand is displayed in Figure 1.
2.1.2 Range of Motion
Cobos et al. [15, Table III] provide a concise overview of the range of motion for each joint. [^1] Range of flexion seems to be swapped between MCP and PIP joints [41]. However, the exact range depends on the person. For example, a human can have distal hyperextensibility of the thumb (hitchhiker’s thumb), which allows them to extend their thumb’s IP joint by 90 degrees so that its active range of motion is 180 degrees. The Kapandji index [40] measures the range of motion of opposition/reposition in the thumb’s TMC joint to evaluate an individual’s dexterity. Furthermore, there is a range that can be actively reached through muscle contraction and it is possible to extend movement in some joints even further through contact with the environment. For example, passive extension in the DIP joints can be about 30 degrees [41], which is considerably larger than the active extension of 5 degrees.
@@ 2.1.3 Muscles and Tendons
Figure 3: Muscles of the human hand. ByYoung et al. [118], $C$ BY 4.0.
@@
Several muscles work together to flex and extend the wrist joint. Flexor carpi radialis, flexor carpi ulnaris, and palmaris longus flex the wrist joint. Extensor carpi ulnaris, extensor carpi radialis longus, and extensor carpi radialis brevis extend the wrist joint [118]. However, most of these muscles also have another function. All carpi radialis muscles contribute to abduction, and all carpi ulnaris muscles contribute to adduction of the wrist. Furthermore, there are extrinsic finger muscles that contribute to wrist flexion and extension [118].
@@
The thumb is an unusual yet essential finger with its large range of motion in the TMC joint and ability to oppose the other fingers. Eight muscles control it. The two degrees of freedom of the TMC joint are controlled by all intrinsic (opponens pollicis, abductor pollicis brevis, flexor pollicis brevis, adductor pollicis) and extrinsic (flexor pollicis longus, extensor pollicis longus, extensor pollicis brevis, abductor pollicis longus) muscles of the thumb. The MCP joint is flexed by flexor pollicis longus and brevis and extended by extensor pollicis longus and brevis. The IP joint is only flexed by flexor pollicis longus and extended by extensor pollicis longus. Abductor pollicis longus also contributes to wrist abduction, and flexor pollicis longus to wrist flexion.
@@
2.1.4 Redundancy
Often, multiple muscles work together to perform one movement. For example, seven muscles contribute to controlling the two degrees of freedom of the wrist. This, however, does not mean that muscles are redundant and can be removed, as [48] point out.
@@ 2.1.5 Force
[113] report maximal power grip forces of 450 N. [111] observe up to 662 N for the 90th percentile of 25–29 year old men in their dominant hand.
2.1.6 Coupling
It is not possible to completely isolate movements of each degree of freedom. Flexion of one finger causes nearby fingers to move as well (enslaving force phenomenon [93]). The effect increases with the force applied to the main finger. Mechanical coupling between tendons and muscle compartments, multi-digit motor units that co-activate, and overlapping representations in the brain are possible causes. For example, FDP flexes all three finger joints [4].
Another example of coupling is abduction of the index finger, which is facilitated by the dorsal interossei muscle, thereby pulling the thumb closer to the index finger. This movement of the thumb must be compensated for by other muscles, resulting in small movements. However, [33] show that the thumb is the finger that moves most independently. An independently movable and opposable thumb is essential for the prehensile function of the hand.
2.1.7 Variations
The exact dimensions and the range of motion vary from person to person. Even the number of fingers can be different, as some people might have six fingers [59].
@@ 2.2.1 Biomechanical aspects of haptic perception
Haptic perception is often linked to the sense of touch, mediated primarily by skin receptors that fall into three categories based on their function: mechanoreceptors for pressure and vibration, thermoceptors for temperature changes, and nociceptors for pain [78, Chap. 9]. These receptors allow us to perceive size, shape, texture, and temperature, which are crucial for object manipulation [18].
Proprioception, or the sense of self-movement and body position, is integral to haptic perception. This sense originates from receptors in muscles, joints, and tendons [53, 17], playing a key role in perceiving object properties like shape through the alignment of bones and muscle stretching [70]. Proprioceptive receptors in muscles, skin, and joint capsules [97] are crucially connected with motor control [100].
Muscle spindles sense changes in muscle length, while Golgi tendon organs (GTOs) measure muscle-applied force [97], contributing to detailed representations of hand and limb positions. Joint capsule receptors measure extreme positions of the joints to prevent overextension [97]. The proprioceptive signals are highly interconnected with motor control [100] and have a fast conduction velocity of 80–120 m/s [92]. These proprioceptive signals, interacting with tactile feedback, convey information about the state of hand, wrist, and forearm muscles.
@@
Cutaneous nerve fibers on the dorsal hand surface react to skin stretch during joint movements and convey joint angle information [57]. Stimulating the skin on the dorsal surface can induce illusions of movement. However, this effect has only been tested where a single joint was deflected [57].
The brain devotes a significant portion of the somatosensory cortex to the hand, reflecting its behavioral importance. About 20% of the surface in Brodmann’s areas [57] are allocated to hand representations.
2.2.2 Skills enabled by haptic perception
Haptic perception enables humans to explore objects using strategies known as exploratory procedures (EPs) [52, 53]. These EPs are categorized based on their focus: the substance of an object (texture, hardness, temperature, weight), structural properties (global shape, volume, weight), and function discovery (movable parts, potential functionality). For instance, the lateral motion EP moves skin over an object’s surface to assess texture, while the pressure EP tests hardness through poking or tapping. Static contact EP assesses temperature briefly through passive touch. Other EPs include unsupported holding for weight inference, enclosing for global shape, contour following for detail, and part motion testing and function testing for understanding movement and function. Figure 5 shows examples of the exploratory procedures for the first two categories.
@@ 2.2.2 Skills enabled by haptic perception
Proprioception, which is the perception of self-movement and body position, plays a crucial role in daily activities and fine motor skills [57]. Lack of proprioception can lead to unstable reach trajectories, impaired postural stability, and poor hand coordination, even when visual feedback is available [53]. Stereognosis, the haptic perception of three-dimensional shapes, requires integrating proprioceptive with cutaneous signals [57]. Proprioceptive signals describe hand conformation during grasp, while tactile signals deliver contact location and geometric features such as edges and corners at each contact point [57]. This integration ensures effective object perception and showcases how multimodal signals encode object identity even before contact occurs [57].
@@ 2.2.3 Role in Grasping
Effective grasping demands fine control over finger muscle strength and precise timing throughout various phases of grasp. A lack of tactile sensing can prolong the duration of the finger opening phase, impairing grasp control and coordination [18]. Tactile cues at the beginning and end of movements enable the optimization of grasp temporal parameters, facilitating an understanding the timing of sequential actions, such as playing an instrument or executing synchronized movements [18].
While tactile senses are pivotal, the human sensory system is inherently multimodal, integrating touch, vision, and auditory feedback to construct a comprehensive internal representation known as the “body schema” [18]. This integration increases the reliability of sensory estimates and enhances robustness [18]. In grasping, a precise measure of contact forces is essential for force control and maintaining stable grasps [18].
In addition to understanding the magnitude of forces, the direction is critical for dexterous manipulation, as it regulates the balance between normal and tangential forces to ensure grasp stability, known as the “friction cone” [18]. Shear forces provide crucial insights into shape, surface texture, and slip, forming a significant part of the tactile information required for interacting with objects and discerning material properties such as hardness and temperature [18].
@@ 2.3.1 Central Nervous System (CNS)
We summarize the role of the CNS for motor control based on [10].
@@ 2.3.2 Motor Program Theory
Motor program theory explains learning, representation, and execution of fast, open-loop motor skills that are realized in the pontocerebellum. It is based on the observations that reaction time increases with complexity of movements [76], animals that are deprived of sensory information using a surgical process of deafferentation still produce effective movements [76], and muscular activity patterns during the first ca. 100 ms of limb movement remain the same even when a person’s limb is unexpectedly prevented from moving [109]. [86, 87] proposed generalized motor programs (GMPs) and the schema theory for learning fast, open-loop motor skills. GMPs are representations of complex, coordinated, and rapid movements that are learned and allow modification along several dimensions, such as movement time, movement amplitude, and the effector system used to produce the action, thereby preserving the temporal pattern of the movement. Schema theory extends GMPs. It suggests that tuples of situation / context parameters, GMP parameters, extrinsic feedback (result of movement), and internal feedback (sensor information) are used to form schemas. When encountering a situation, an appropriate GMP and parameterization are selected to achieve the desired result.
@@ 2.3.3 Dynamical Systems Theory
Another theory is needed to explain closed-loop, sensorimotor control. An interesting observation is that humans have many degrees of freedom, e.g., in their hands and arms, that need to be coordinated, many more degrees of freedom than are necessary to solve specific tasks. [6] found that humans reliably reach high-level goals despite high variability in their movements. A related phenomenon is the so-called uncontrolled manifold [88], which states that several degrees of freedom are actively controlled and stable. In contrast, others are not controlled and show high variability. Dynamical systems theories [94] explain these observations through concepts such as muscle synergies and coordinative structures that result in attractor states or dynamic stability with many redundant degrees of freedom [89, 28]. The concept of synergies has been applied to hand control by [58], who demonstrate that reach-to-grasp motions can be produced with a base posture with refinements of finger positions. The first three dimensions of a singular value decomposition account for 99.5% of the variance of all observed grasps. Hence, it is assumed that actively controlled degrees of freedom are considerably lower than the actual number of degrees of freedom of the hand. [84] apply the concept to forces. Synergies describe the outcome, but their origin needs an additional explanation.
@@ 2.3.4 Stochastic Optimal Feedback Control
[102, 101] propose stochastic optimal feedback control to combine performance guarantees inherent in optimization models with behavioral richness emerging from dynamical systems. [102] argue that planning and execution are not separated. This framework explains task-constrained variability, goal-directed corrections, motor synergies, and controlled parameters, simplifying rules and discrete coordination modes into a unifying theory of sensorimotor skills.

### v2

@@ Abstract
Human manipulation skills represent a pinnacle of their voluntary motor functions, requiring the coordination of many degrees of freedom and processing of high-dimensional sensor input to achieve such a high level of dexterity. Thus, we attempt to answer whether the human hand, with its associated biomechanical properties, sensors, and control mechanisms, is an ideal that we should strive for in robotics—do we really need anthropomorphic robotic hands?
@@
Although complex five-fingered hands are often used as the ultimate goal for robotic manipulators, they are not necessary for all tasks. We found that wrist flexibility and finger abduction/adduction are often more important for manipulation capabilities. Increasing the number of fingers, actuators, or degrees of freedom is not always necessary. Three fingers often are a good compromise between simplicity and dexterity. Non-anthropomorphic hand designs with two opposing pairs of fingers or human hands with six fingers can further increase dexterity, suggesting that the human hand is not the optimum. Consequently, we argue for function-based rather than form-based biomimicry.
@@
Robotics Hands , Grippers , Smart Prosthetics , Manipulation
[…]
[1]organization=German Research Center for Artificial Intelligence GmbH (DFKI), Robotics Innovation Center, addressline=Robert-Hooke-Straße 1, postcode=28359, city=Bremen, country=Germany
[…]
[2]organization=Leibniz Universität Hannover, L3S Research Center, addressline=Appelstraße, postcode=30167, city=Hanover, country=Germany
@@ 1 Introduction
All of which prompts the question of whether we need anthropomorphic robotic hands. Attempting to answer this question, we analyze the human hand and its capabilities, including kinematics and perception, to find characteristic features of anthropomorphic hands. Furthermore, we review commercially available robotic and prosthetic hands to identify anthropomorphic features that are technically feasible. Then we perform a systematic literature review to correlate hand features derived from the previous two sections, skill categories, and manipulable degrees of freedom. We discuss and connect our findings to literature from neuroscience, biology, prosthetics, and robotics.
@@ 2 The Human Hand
2.1 Biomechanics
Figure 1: Kinematics of the human hand (legend in Figure 7(a)).
The human hand is one of the most advanced manipulation devices that can be found in nature [5]. We summarize common characteristics of most human hands. The exact dimensions and the range of motion vary from person to person. Even the number of fingers can be different, as some people might have six fully functional fingers [3].
@@ 2.1.1 Joints and Degrees of Freedom
The human hand has many bones (see Figure 2), but not all of them are connected through movable joints. It has at least 20 degrees of freedom—four in each finger. Index, middle, ring, and little finger can perform flexion and extension in three joints (metacarpo-phalangeal joint, MCP; proximal interphalangeal joint, PIP; distal interphalangeal joint, DIP) and adduction (moving closer to the middle finger) and abduction (moving away from the middle finger) in one joint (MCP). The thumb is a special case with a saddle joint (trapezo-metacarpal joint, TMC) that has two degrees of freedom to achieve flexion/extension and opposition/reposition, and two more joints (MCP and interphalangeal joint, IP) to perform flexion/extension. It is possible to move the metacarpal bones and hollow the palm to a limited extent [5] so that Cobos at el. [7] define hand kinematics with 24 degree of freedom. In addition, assuming the carpus to be rigid is only an approximation [5], and the palm, as well as the skin, are soft and deformable. A kinematic diagram of the hand is displayed in Figure 1.
Cobos at el. [7, Table III] provide a concise overview of the range of motion for each joint. [^1] Range of flexion seems to be swapped between MCP and PIP joints [5]. However, the exact range varies from person to person. For example, a human can have distal hyperextensibility of the thumb (hitchhiker’s thumb), which allows them to extend their thumb’s IP joint by 90 degrees so that its active range of motion is 180 degrees. Furthermore, there is a range that can be actively reached through muscle contraction and it is possible to extend movement in some joints even further through contact with the environment. For example, passive extension in the DIP joints can be about 30 degrees [5], which is considerably larger than the active extension of 5 degrees.
@@ 2.1.2 Muscles and Tendons
Figure 3: Muscles of the human hand. By [6], $C$ BY 4.0.
@@
Several muscles work together to flex and extend the wrist joint. [^2] Flexor carpi radialis, flexor carpi ulnaris, and palmaris longus flex the wrist joint. Extensor carpi ulnaris, extensor carpi radialis longus, and extensor carpi radialis brevis extend the wrist joint [6]. However, most of these muscles also have another function. All carpi radialis muscles contribute to abduction, and all carpi ulnaris muscles contribute to adduction of the wrist. Furthermore, there are extrinsic finger muscles that contribute to wrist flexion and extension [6].
@@
The thumb is an unusual yet essential finger with its large range of motion in the TMC joint and ability to oppose the other fingers. Eight muscles control it. The two degrees of freedom of the TMC joint are controlled by all intrinsic [^3] Intrinsic muscles that control the thumb: opponens pollicis, abductor pollicis brevis, flexor pollicis brevis, adductor pollicis. and extrinsic [^4] Extrinsic muscles that control the thumb: flexor pollicis longus, extensor pollicis longus, extensor pollicis brevis, abductor pollicis longus. muscles of the thumb. The MCP joint is flexed by flexor pollicis longus and brevis and extended by extensor pollicis longus and brevis. The IP joint is only flexed by flexor pollicis longus and extended by extensor pollicis longus. Abductor pollicis longus also contributes to wrist abduction, and flexor pollicis longus to wrist flexion.
@@
2.1.3 Lack of Muscle Redundancy and Coupling of Degrees of Freedom
Often, multiple muscles work together to perform one movement. For example, seven muscles contribute to controlling the two degrees of freedom of the wrist. This, however, does not mean that muscles are redundant and can be removed. For instance, Kutch and Valero-Cuevas [10] point out that cocontraction of specific muscles is often required to produce certain forces with the index finger and that deactivating individual muscles can considerably impair motor function. In their experiments, less than 5% of feasible forces are robust to loss of any one muscle.
It is also not possible to completely isolate movements of each degree of freedom. Flexion of one finger causes nearby fingers to move as well (enslaving force phenomenon [11]). The effect increases with the force applied to the main finger. Mechanical coupling between tendons and muscle compartments, multi-digit motor units that co-activate, and overlapping representations in the brain are possible causes. For example, FDP flexes all three finger joints [12].
Another example of coupling is abduction of the index finger, which is facilitated by the dorsal interossei muscle, thereby pulling the thumb closer to the index finger. This movement of the thumb must be compensated for by other muscles, resulting in small movements. However, Ingram et al. [13] show that the thumb is the finger that moves most independently. An independently movable and opposable thumb is essential for the prehensile function of the hand.
@@ 2.1.4 Force
The human hand has a high force-to-weight ratio. For example, Wang et al. [14] observe grip forces of 662 N for the strongest analyzed group (25–29 year old men in their dominant hand, 90th percentile). They measure grip forces with a dynamometer. All fingers of the human hand work together to produce the maximum force that is possible.
@@ 2.2.1 Biomechanical Aspects of Haptic Perception
Haptic perception is often linked to the sense of touch, mediated primarily by skin receptors that fall into three categories based on their function: mechanoreceptors for pressure and vibration, thermoceptors for temperature changes, and nociceptors for pain [16, Chap. 9]. These receptors allow us to perceive size, shape, texture, and temperature, which are crucial for object manipulation [17].
@@
Proprioception, or the sense of self-movement and body position, is integral to haptic perception. This sense originates from receptors in muscles, joints, and tendons [19, 20], playing a key role in perceiving object properties, such as shape, through the alignment of bones and muscle stretching [21]. Proprioceptive receptors in muscles, skin, and joint capsules [22] are crucially connected with motor control [23], and have a fast conduction velocity of 80–120 m/s [24]. These proprioceptive signals, interacting with tactile feedback, convey information about the state of hand, wrist, and forearm muscles.
Muscle spindles sense changes in muscle length, while Golgi tendon organs (GTOs) measure muscle-applied force [22], contributing to detailed representations of hand and limb positions. Joint capsule receptors measure extreme positions of the joints to prevent overextension [22].
Cutaneous nerve fibers on the dorsal hand surface react to skin stretch during joint movements and convey joint angle information [25]. Stimulating the skin on the dorsal surface can induce illusions of movement. However, this effect has been tested only when a single joint was deflected [25].
2.2.2 Hand Representation and Object Perception
The brain devotes a significant portion of the somatosensory cortex to the hand, reflecting its behavioral importance. About 20% of the surface in Brodmann’s areas [25] are allocated to hand representations. Stereognosis, the haptic perception of three-dimensional shapes, requires integrating proprioceptive with cutaneous signals [25]. Proprioceptive signals describe hand conformation during grasp, while tactile signals deliver contact location and geometric features such as edges and corners at each contact point [25]. This integration ensures effective object perception and showcases how multimodal signals encode object identity even before contact occurs [25].
2.2.3 Object Exploration
@@
Haptic perception enables humans to explore objects using strategies known as exploratory procedures (EPs) [27, 19]. These EPs are categorized based on their focus: the substance of an object (texture, hardness, temperature, weight), structural properties (global shape, volume, weight), and function discovery (movable parts, potential functionality). For instance, the lateral motion EP moves skin over an object’s surface to assess texture, while the pressure EP tests hardness through poking or tapping. Static contact EP assesses temperature through passive touch. Other EPs include unsupported holding for weight inference, enclosing for global shape, contour-following for detail, and part-motion testing and function testing to understand movement and function. Figure 5 shows examples of the exploratory procedures for the first two categories.
@@ 2.2.4 Role in Grasping
Effective grasping demands fine control of finger muscle strength and precise timing across the various phases of the grasp. A lack of tactile sensing can prolong the finger-opening phase, impairing grasp control and coordination [17]. Tactile cues at the beginning and end of movements enable the optimization of grasp temporal parameters, facilitating an understanding of the timing of sequential actions, such as playing an instrument or executing synchronized movements [17].
In addition to understanding the magnitude of forces, the direction is critical for dexterous manipulation, as it regulates the balance between normal and tangential forces to ensure grasp stability, known as the “friction cone” [17]. Shear forces provide crucial insights into shape, surface texture, and slip, and they form a significant part of the tactile information required for interacting with objects and discerning material properties such as hardness and temperature [17].
@@ 2.3.1 Central Nervous System (CNS)
We summarize the role of the CNS for motor control based on Brandes et al. [28].
@@ 2.3.2 Motor Program Theory
Motor program theory explains learning, representation, and execution of fast, open-loop motor skills that are realized in the pontocerebellum. It is based on the observations that reaction time increases with complexity of movements [29], animals that are deprived of sensory information using a surgical process of deafferentation still produce effective movements [29], and muscular activity patterns during the first ca. 100 ms of limb movement remain the same even when a person’s limb is unexpectedly prevented from moving [30]. Schmidt [31, 32] proposed generalized motor programs (GMPs) and the schema theory for learning fast, open-loop motor skills. GMPs are representations of complex, coordinated, and rapid movements that are learned and allow modification along several dimensions, such as movement time, movement amplitude, and the effector system used to produce the action, thereby preserving the temporal pattern of the movement. Schema theory extends GMPs. It suggests that tuples of situation / context parameters, GMP parameters, extrinsic feedback (result of movement), and internal feedback (sensor information) are used to form schemas. When encountering a situation, an appropriate GMP and parameterization are selected to achieve the desired result.
@@ 2.3.3 Dynamical Systems Theory
Another theory is needed to explain closed-loop, sensorimotor control. An interesting observation is that humans have many degrees of freedom, e.g., in their hands and arms, that need to be coordinated, many more degrees of freedom than are necessary to solve specific tasks. Bernstein [33] found that humans reliably reach high-level goals despite high variability in their movements. A related phenomenon is the so-called uncontrolled manifold [34], which states that several degrees of freedom are actively controlled and stable. In contrast, others are not controlled and show high variability. Dynamical systems theories [35] explain these observations through concepts such as muscle synergies and coordinative structures that result in attractor states or dynamic stability with many redundant degrees of freedom [36, 37]. The concept of synergies has been applied to hand control by Mason et al. [38], who demonstrate that reach-to-grasp motions can be produced with a base posture with refinements of finger positions. The first three dimensions of a singular value decomposition account for 99.5% of the variance of all observed grasps. Hence, it is assumed that actively controlled degrees of freedom are considerably lower than the actual number of degrees of freedom of the hand. Santello and Soechting [39] apply the concept to forces. Synergies describe the outcome, but their origin needs an additional explanation.
@@ 2.3.4 Stochastic Optimal Feedback Control
Todorov et al. [40, 41] propose stochastic optimal feedback control to combine performance guarantees inherent in optimization models with behavioral richness emerging from dynamical systems. Todorov and Jordan [40] argue that planning and execution are not separated. This framework explains task-constrained variability, goal-directed corrections, motor synergies, and controlled parameters, simplifying rules and discrete coordination modes into a unifying theory of sensorimotor skills.